Blocks A and B have weights of 44 N and 22 N, respectively. (a) Determine the minimum weight of block C to keep A from sliding if between A and the table is 0.20. (b) Block C suddenly is lifted off A. What is the acceleration of block A if 3 between A and the table is 0.15?


1
Expert's answer
2021-02-22T10:22:47-0500

a)

T=22 N,

"N=\frac{T}{\mu _s}=110~N,"

W=110-44=66 N,

b)

"\begin{cases}
 T-\mu_k N=m_A a, \\
 m_B g-T=m_B a;
\end{cases}"

"a=\frac{9.8\cdot (22-0.15\cdot 44)}{66}=2.3 ~\frac{m}{s^2}."